A bridge is 176.4 m above a river. If a lead-weighted fishing line is thrown from the bridge with a horizontal velocity of 22.0 m/s, how far has it moved horizontally when it hits the water?
1
Expert's answer
2012-10-15T12:14:01-0400
The falling time is

T = √(2H/g) = √(2*176.4[m]/9.8[m/s²]) = 6.0[s].

Therefore, fishing line will move

D = 22.0[m/s]*6.0[s] = 132[m]

horizontally when it hits the water.